# Patronus: Bringing Transparency to Diffusion Models with Prototypes

### v1

@@ Abstract
Diffusion-based generative models, such as Denoising Diffusion Probabilistic Models (DDPMs), have achieved remarkable success in image generation, but their step-by-step denoising process remains opaque, leaving critical aspects of the generation mechanism unexplained. To address this, we introduce Patronus, an interpretable diffusion model inspired by ProtoPNet. Patronus integrates a prototypical network into DDPMs, enabling the extraction of prototypes and conditioning of the generation process on their prototype activation vector. This design enhances interpretability by showing the learned prototypes and how they influence the generation process. Additionally, the model supports downstream tasks like image manipulation, enabling more transparent and controlled modifications. Moreover, Patronus could reveal shortcut learning in the generation process by detecting unwanted correlations between learned prototypes. Notably, Patronus operates entirely without any annotations or text prompts. This work opens new avenues for understanding and controlling diffusion models through prototype-based interpretability. Our code is available at https://github.com/nina-weng/patronus.
@@ 1 Introduction
The generative capabilities of modern machine learning models, particularly diffusion models, have advanced significantly, enabling the creation of highly realistic samples that closely resemble real-world data. However, their opacity raises critical concerns, including bias amplification [22], unsafe content [27], and copyright violations [31]. Their lack of transparency makes it difficult to detect and mitigate these risks, highlighting a fundamental question: How can the diffusion generation process be interpreted and understood? Specifically, what visual patterns emerge during generation, where and when they appear, and to what extent they can be controlled. Addressing these questions is essential – not only for improving generative modeling but also for ensuring interpretability, transparency, and ethical deployment, aligning with regulatory frameworks such as the EU AI Act.
Figure 1: Interpretability, manipulation, and diagnostic capabilities of Patronus. Interpretability: By integrating a prototypical network as the encoder, Patronus learns semantic prototypes (“what”) and explains the generative process by revealing where and when they emerge. Manipulation: By adjusting the prototype activation vector, Patronus enables targeted semantic image editing. Diagnosis: Semantic image editing may reveal unwanted correlations learned from the training data during generation. Note that Patronus operates without any annotations or text prompts.
Existing approaches to improving interpretability in diffusion-based visual generation typically fall into two categories. The first relies on post-hoc analysis to investigate how semantic information are encoded in intermediate representations [16, 18, 25, 10, 29, 30]. However, this method is inherently retrospective and limited in its ability to provide direct control over generation. The second approach introduces additional encoder-based semantic vectors for diffusion guidance [26, 19, 33], which improves controllability, but often resulting in representations that are difficult to interpret. Moreover, these methods tend to capture global (e.g. face shape, pose) rather than fine-grained patterns (e.g. hair details, facial expressions, make-up details), and the latter are crucial for interpretability.
@@
Drawing inspiration from ProtoPNet [4], a deep learning framework that enhances interpretability through part-prototypical networks, we propose Patronus, a novel diffusion-based generative model augmented with a prototypical network for semantic encoding. Our approach trains prototype features to capture localized patterns within image patches, and uses prototype activation vector to encode the presence of semantic information for diffusion guidance. This design effectively reduces the latent dimensionality while preserving sufficient semantic information. Furthermore, by manipulating the activation vector as conditioning signals, we enable direct visualization of the learned prototypes, enhancing interpretability and control.
@@
  1. 1.
    We propose Patronus: An interpretable diffusion model for image generation which incorporates a prototypical network for prototype learning and representation, alongside a conditional guidance by a prototype activation vector —entirely without extra annotations.
  2. 2.
    We introduce a novel method for visualizing learned prototypes, revealing how each prototype influences the generation process, and enabling manipulation of the prototype activation vector for image generation.
  3. 3.
@@
  4. 4.
    We demonstrate the potential of Patronus to diagnose unwanted correlations in datasets, offering a valuable tool for mitigating biases in generative models and promoting fairness in their deployment.
@@ 2 Related Work
2.1 Visual Prototype Explanations
Our work is closely related to and motivated by ProtoPNet [4] and its subsequent developments [5, 32, 24, 7], which build interpretable deep learning frameworks by learning prototypes, intermediate representations of visually similar patterns between training and inference images.
Under such a design, a key challenge is visualizing the learned prototypes. Li et al. [20] used a decoder to reconstruct prototypes, while ProtoPNet replaced it by identifying the closest matching encoded representations from the training set via distance comparisons, with the improvement of flexible patch sizes. Follow-up works explore alternative prototype design but do not enhance visualization.
@@ 2.2 Interpretable Diffusion Models
Recent work on explaining diffusion models has focused on the following two directions: (1) Semantic interpretation of the internal features: Investigating how diffusion models encode and represent semantic information within their intermediate layers; and (2) Construction of semantic vectors from an additional encoder for diffusion guidance: Branching out another auto-encoder to construct a semantic vector for diffusion process guidance.
@@
Diffusion models were traditionally viewed as lacking internal representations compared to other generative models, like VAE [14] and GAN [8]. However, Kwon et al. [16] pointed out that diffusion models possess a semantic latent space within the U-Net’s intermediate layer. They leverage the bottleneck for semantic control over the denoising process and propose the asymmetric reverse process to edit the image based on the discovered semantic meaning.
Building on this idea, researchers have sought to interpret the latent space (called $h$-space in [16]) of diffusion models by grouping the latent vectors with different noise schedules [18], applying pullback metrics to obtain meaningful local latent basis [25], and uncovering semantically meaningful directions by both PCA and linear properties of the semantic latent space [10]. Other works [29, 30] extract structural information from the latent space and use it for I2I translation or prompted generation tasks.
@@
An alternative way to enhance the interpretability of diffusion models is to integrate an additional encoder to extract semantic features for guidance. DiffAE [26] pioneered this idea using a learnable encoder for latent semantics. Expanding on this, DiffuseGAE [19] proposed a Group-supervised AutoEncoder module to achieve better latent disentanglement, while InfoDiffusion [33] reduced latent dimensionality and enforced mutual information constraints for more effective learning.
@@
3 Method: Patronus Framework for Interpretable DDPM-based Image Generation
Figure 2: Overview of Patronus: contains a prototypical network for prototype extraction and a conditional DDPM for generation.
Our proposed model, Patronus: Prototype-Assisted Transparent Diffusion Model, is illustrated in Fig. 2. Designed to enhance transparency and interpretability in diffusion models, Patronus incorporates a prototype extraction and representation module (bottom part of Fig. 2). This module learns patch-based prototypes within the image and computes similarity score for each prototype, which are then used to condition the diffusion process (top part of Fig. 2). We elaborate on the details of the prototype extraction and representation module in Sec. 3.1 and the conditional DDPM in Sec. 3.2. Furthermore, we explain how transparency and interpretability are achieved, including sampling strategy and manipulations, in Sec. 3.3. We detail the unconditional sampling strategies in Sec. 3.4, and furthermore prove that adding conditions does not degrade the denoiser training in Sec. 3.5.
@@ 3.1 Prototype Extraction and Representation
This module, inspired by ProtoPNet [4], consists of two key components: The prototype encoder transforms input images into patch-based feature representations. Utilizing the properties of convolutional neural networks (CNNs), each output neuron in the feature map corresponds to a specific patch of the input image, determined by the network’s receptive field. This patch-based representation enables the model to focus on localized patterns and learn fine-grained prototypes. The activation vector is derived by calculating similarity scores for each learned prototype, based on the distance between encoded patches and prototypes, where higher scores indicate stronger matches.
This module works as follows: As shown in Fig. 2, given an input image $x_{0}$, the prototype encoder $f$ extracts features $z=f(x_{0})$ into a tensor of shape $H\times W\times D$. The network learns $m$ prototypes in the latent feature space during training, denoted as $P=\{p_{j}\}_{j=1}^{m}$, each with the shape $1\times 1\times D$111whose generalization to $H_{1}\times W_{1}\times D$ as in [4] is straightforward.. Each prototype $p_{j}$ can be interpreted as a latent encoding of a patch in the original pixel space. This patch, importantly, need not exist in the dataset but should lie within the plausible data distribution.
@@
The sequence of minimum distances for all prototypes is converted into an activation vector $s$ using a log transformation: $s=\log(\frac{d^{2}+1}{d^{2}+\epsilon})$ , where $\epsilon$ is a small positive constant.
@@ 3.2 Conditional Diffusion Process
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (3)
@@
$p_{\theta}(x_{t-1}\mid x_{t},s)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,s),\Sigma_{\theta}(x_{t},t))$ (4)
@@
$\mathcal{L}_{ddpm}=\mathbb{E}_{x_{0},\theta,t}[||\epsilon-\epsilon_{\theta}(x_{t},t,s)||^{2}],$ (5)
@@ 3.3 Transparency and Interpretability of Patronus
Integrating a prototypical network as a semantic meaning extraction module brings inherent interpretability: Each learned prototype vector $p_{j}$ represents a patch in the image domain. In ProtoPNet [4], those patches are retrieved via a greedy search of the training set and finding the patch with closest embedding to $p_{j}$.
@@
Note that this method could also be used to visualize prototypes in other prototype based deep learning models.
@@
Both the visualization of prototypes and their manipulation via activation vectors build on the DDPM sampling process. However, for stricter control over the stochasticity introduced by random noise, the Denoising Diffusion Implicit Models (DDIM) sampler provides an alternative approach:
$p_{\theta}(x_{t-1}\mid x_{t},s)=\mathcal{N}(x_{t-1};\mu_{t},\sigma^{2}\mathbf{I})$ (7)
$\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}}\cdot{\epsilon}_{\theta}(x_{t},t,s)$ (8)
$\sigma^{2}=\eta^{2}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot(1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}})$ (9)
By setting $\eta=0.0$, the process becomes deterministic, leading to the following update for the reverse process:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot{\epsilon}_{\theta}(x_{t},t,s),$ (10)
where $\hat{x}_{0}$ is the estimated denoised image, computed as:
$\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot{\epsilon}_{\theta}(x_{t},t,s)).$ (11)
This DDIM sampling requires $x_{T}$, which represents the initial noise in the diffusion process. This could be obtained by performing a deterministic backward generative process:
$x_{t+1}=\sqrt{\bar{\alpha}_{t+1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t+1}}\cdot{\epsilon}_{\theta}(x_{t},t,s)$ (12)
@@ 3.4 Unconditional Sampling Strategy
For unconditional sampling, we adopt the approach of [26], training an auxiliary latent diffusion model $p(s_{t-1}|s_{t},t)$ to sample $s$. During training, we first jointly optimize the prototypical encoder with the conditional DDPM; subsequently, the latent diffusion model is trained while keeping the previously learned parameters fixed.
@@ 3.5 Adding the Condition to the Objective
In Patronus, the prototype encoder is jointly optimized with the denoiser. To show that this simultaneous training does not degrade the generated distribution, we analyze how updating the condition $s$ affects the likelihood. For ease of derivation, we use the Evidence Lower Bound (ELBO) as an equivalent objective to generalize denoising losses [11].
@@
Optimizing the conditioning variable $s$ with denoising loss cannot degrade the quality of the results.
@@
Assume that $s^{i}$ is updated to $s^{i+1}$via a gradient step optimizing the ELBO. By definition of gradient descent:
$\text{ELBO}(s^{i})\leq\text{ELBO}(s^{i+1})$
The ELBO loss is given by:
$\text{ELBO}(s^{i})=\mathbb{E}\left[\log p(x_{0},x_{1:T}|s^{i})-\log q(x_{1:T}|x_{0})\right]$
Since the ELBO is a variational lower bound on the marginal log-likelihood $\log p(x_{0})$, an increase in ELBO generally implies that the approximate posterior distribution $q(x_{1:T}|x_{0})$ better captures the true posterior. However, increasing ELBO does not necessarily increase the likelihood $p(x_{0}|s)$ itself. We therefore consider two possible cases:
Case 1:
If $p(x_{0}|s^{i})\leq p(x_{0}|s^{i+1}),$ then the update improves the likelihood of $x_{0}$ under the model, directly enhancing the sample quality, as higher likelihood generally corresponds to more accurate or better-calibrated samples. Here, optimization has both increased the ELBO and improved the likelihood.
Case 2:
If instead, we have $p(x_{0}|s^{i})>p(x_{0}|s^{i+1})>p(x_{0}),$ then the likelihood has not increased, but it remains greater than the marginal likelihood $p(x_{0})$. In this case, we analyze the difference between likelihood and ELBO:
$p(x_{0}|s^{i})-\text{ELBO}(s^{i})>p(x_{0}|s^{i+1})-\text{ELBO}(s^{i+1}).$
As the gap between likelihood and ELBO shrinks, ELBO becomes a tighter approximation to the true marginal log-likelihood $\log p(x_{0})$. Thus, while the likelihood itself may not have increased, our approximation is now more reliable.
In both cases, optimizing $s$ has positive impact and cannot degrade the quality of the results: Improved likelihood and better sample quality (Case 1), or a tighter ELBO bound, refining likelihood approximation (Case 2). ∎
This analysis is not applicable where a different objective is used to train the latent condition (e.g. InfoDiffusion [33]).

### v2

# Patronus: Interpretable Diffusion Models with Prototypes

@@ Abstract
Uncovering the opacity of diffusion-based generative models is urgently needed, as their applications continue to expand while their underlying procedures largely remain a black box. With a critical question – how can the diffusion generation process be interpreted and understood? – we proposed Patronus, an interpretable diffusion model that incorporates a prototypical network to encode semantics in visual patches, revealing what visual patterns are modeled and where and when they emerge throughout denoising. This interpretability of Patronus provides deeper insights into the generative mechanism, enabling the detection of shortcut learning via unwanted correlations and the tracing of semantic emergence across timesteps. We evaluate Patronus on four natural image datasets and one medical imaging dataset, demonstrating both faithful interpretability and strong generative performance. With this work, we open new avenues for understanding and steering diffusion models through prototype-based interpretability.
Our code is available at https://github.com/nina-weng/patronus.
@@ 1 Introduction
The generative capabilities of modern machine learning models, particularly diffusion models, have advanced significantly, enabling the creation of highly realistic samples that closely resemble real-world data. However, their opacity raises critical concerns, including bias amplification (luccioni2023stable), unsafe content (qu2023unsafe), and copyright violations (vyas2023provable). Their lack of transparency makes it difficult to detect and mitigate these risks, highlighting a fundamental question: How can the diffusion generation process be interpreted and understood? Specifically, what visual patterns emerge during generation, where and when they appear, and to what extent they can be controlled. Addressing these questions is essential, not only for improving generative ability but also for ensuring interpretability, transparency, and ethical deployment, aligning with regulatory frameworks such as the EU AI Act.
Existing approaches to improving interpretability in diffusion-based visual generation typically fall into two categories. The first relies on post-hoc analysis to investigate how semantic information are encoded in intermediate representations (kwon2022diffusion; lee2023diffusion; park2023understanding; haas2024discovering). However, this method is inherently retrospective and limited in its ability to provide direct control over generation. The second approach introduces additional encoder-based semantic vectors for diffusion guidance (preechakul2022diffusion; leng2023diffusegae; wang2023infodiffusion), which improves controllability, but often resulting in representations that are difficult to interpret. Moreover, these methods tend to capture global (e.g. face shape, pose) rather than fine-grained patterns (e.g. hair/make-up details, facial expressions), and the latter are crucial for interpretability.
Figure 1: Proposed Patronus model. (a) Interpretability: By integrating a prototypical network as the encoder, Patronus learns semantic prototypes (“what”) and explains the generative process by revealing where and when they emerge. (b) Diagnosis: Patronus could detect unwanted correlation (e.g., in this case hair color and smile) learned from training data. (c) Overview of Patronus: contains a prototypical network for prototype extraction and a conditional DDPM for generation.
@@
In this work, we propose Patronus, an interpretable diffusion-based generative model integrated with a prototypical network for semantic encoding (Fig. 1). Our approach trains prototype features to capture localized patterns within image patches, and uses prototype activation vector to encode the presence of semantic information for diffusion guidance. This design effectively reduces the latent dimensionality while preserving sufficient semantic information. Patronus allows directly visualizing the semantics of the learned prototypes by manipulating the condition signals on specific prototypes, thereby enhancing interpretability and enabling the diagnosis of shortcut learning. Beyond that, Patronus further reveals when prototypes emerge during the denoising process, providing temporal insights and guiding more efficient editing.
@@
  • We propose Patronus: An interpretable diffusion model for image generation which incorporates a prototypical network for prototype learning and representation, alongside a conditional guidance by a prototype activation vector – entirely without extra annotations.
  • We introduce a novel method for visualizing learned prototypes, revealing the semantic meanings of them and how they engaged in the generation process.
@@
  • We empirically demonstrate the potential of Patronus to diagnose hidden bias by detecting shortcut learning, offering a valuable tool for mitigating biases in generative models and promoting fairness in their deployment.
@@ 2 Related Work
2.1 Prototype-based Interpretability
Our work is closely related to and motivated by ProtoPNet (chen2019looks), which builds interpretable deep learning frameworks by learning prototypes, intermediate representations of visually similar patterns between training and inference images.
Under such a design, a key challenge is visualizing the learned prototypes. li2018deep used a decoder to reconstruct prototypes, which resulted in blurry visualizations due to data ambiguity. ProtoPNet instead identified the closest matching encoded representations from the training set via distance comparisons. However, this approach cannot guarantee capturing the true representative mode of the distribution. Follow-up works (donnelly2022deformable; wang2023learning; ming2019interpretable; ghosal2021multi) explore alternative prototype design but do not enhance visualization. We propose a new method for prototype visualization that allow finding the most probable visual representation of a prototype. Additionally, the aforementioned studies only focus on classification, our work focus on learning and visualizing prototypes for generative models.
@@ 2.2 Interpretable Diffusion Models
Recent work on explaining diffusion models has focused on the following two directions.
@@
Diffusion models were traditionally viewed as lacking internal representations compared to other generative models, like VAE (kingma2013auto) and GAN (goodfellow2014generative). However,  kwon2022diffusion pointed out that diffusion models possess a semantic latent space within the U-Net’s intermediate layer. They leverage the bottleneck for semantic control over the denoising process and propose the asymmetric reverse process to edit the image based on the discovered semantic meaning.
Building on this idea, researchers have sought to interpret the latent space of diffusion models by grouping the latent vectors with different noise schedules (lee2023diffusion), applying pullback metrics to obtain meaningful local latent basis (park2023understanding), and uncovering semantically meaningful directions by both PCA and linear properties of the semantic latent space (haas2024discovering). Other works (si2024freeu; tumanyan2023plug) extract structural information from the latent space and use it for I2I translation or prompted generation tasks.
This line of work focuses on post-hoc analyses of existing diffusion models, whereas our approach provides intrinsic interpretability by design. Moreover, while prior studies often rely on annotated data or external classifiers to uncover the semantic latent space, Patronus learns prototypes in a fully unsupervised manner.
@@
An alternative way to enhance the interpretability of diffusion models is to integrate an additional encoder to extract semantic features for guidance. DiffAE (preechakul2022diffusion) pioneered this idea using a learnable encoder for latent semantics. Expanding on this, DiffuseGAE (leng2023diffusegae) proposed a group-supervised AutoEncoder to achieve better latent disentanglement, while InfoDiffusion (wang2023infodiffusion) reduced latent dimensionality and enforced mutual information constraints for more effective learning.
@@
3 Patronus: Interpretable Diffusion Model
Our proposed model, Patronus: Prototype-Assisted Transparent Diffusion Model, is illustrated in Fig. 1c. Designed to enhance transparency and interpretability in diffusion models, Patronus incorporates a prototype extraction and representation module (Fig. 1c-bottom). This module learns patch-based prototypes within the image and computes similarity score for each prototype, which are then used to condition the diffusion process (Fig. 1c-top). We elaborate on the details of the prototype extraction and representation module in Sec. 3.1 and the conditional DDPM in Sec. 3.2. Furthermore, we explain how transparency and interpretability are achieved, including sampling strategy and manipulations, in Sec. 3.3. We detail the unconditional sampling strategies in Sec. 3.4, and furthermore prove that adding conditions does not degrade the denoiser training in Sec. 3.5.
@@ 3.1 Prototype Extraction and Representation
This module, inspired by ProtoPNet, consists of two key components: The prototype encoder transforms input images into patch-based feature representations. Utilizing the properties of convolutional neural networks (CNNs), each output neuron in the feature map corresponds to a specific patch of the input image, determined by the network’s receptive field. This patch-based representation enables the model to focus on localized patterns and learn fine-grained prototypes. The activation vector is derived by calculating similarity scores for each learned prototype, based on the distance between encoded patches and prototypes, where higher scores indicate stronger matches.
This module works as follows: As shown in Fig. 1c, given an input image $x_{0}$, the prototype encoder $f$ extracts features $z=f(x_{0})$ into a tensor of shape $H\times W\times D$. In this work, the encoder is a 4-layer Conv–ReLU encoder. The network learns $m$ prototypes in the latent feature space during training, denoted as $P=\{p_{j}\}_{j=1}^{m}$, each with the shape $\times 1\times D11$1whose generalization to $H_{1}\times W_{1}\times D$ as in chen2019looks is straightforward.. Each prototype $p_{j}$ can be interpreted as a latent encoding of a patch in the original pixel space. This patch, importantly, need not exist in the dataset but should lie within the plausible data distribution.
@@ 3.2 Conditional Diffusion Process
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)).$ (3)
@@
$p_{\theta}(x_{t-1}\mid x_{t},s)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,s),\Sigma_{\theta}(x_{t},t)).$ (4)
@@
[…]
@@ 3.3 Transparency and Interpretability of Patronus
Integrating a prototypical network as a semantic meaning extraction module brings inherent interpretability: Each learned prototype vector $p_{j}$ represents a patch in the image domain. In ProtoPNet, those patches are retrieved by greedily searching all candidate patches in the training set for the closest embeddings.
@@
This method could also be used to visualize prototypes in other prototypical deep learning models.
@@
Both the visualization of prototypes and their manipulation via activation vectors build on the DDPM sampling process. However, for stricter control over the stochasticity introduced by random noise, the Denoising Diffusion Implicit Models (DDIM) sampler provides an alternative approach. Here, the denoiser is given by $p_{\theta}(x_{t-1}\mid x_{t},s)=\mathcal{N}(x_{t-1};\mu_{t},\sigma^{2}\mathbf{I})$, with mean $\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}}\cdot{\epsilon}_{\theta}(x_{t},t,s)$ and variance $\sigma^{2}=\eta^{2}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot(1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}})$. By setting $\eta=0.0$, the process becomes deterministic, leading to the following update for the reverse process: $x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot{\epsilon}_{\theta}(x_{t},t,s)$. Here, $\hat{x}_{0}$ is the estimated denoised image, computed as $\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot{\epsilon}_{\theta}(x_{t},t,s))$.
This DDIM sampling requires $x_{T}$, which represents the initial noise in the diffusion process. This could be obtained by performing a deterministic backward generative process, given by $x_{t+1}=\sqrt{\bar{\alpha}_{t+1}}\cdot\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t+1}}\cdot{\epsilon}_{\theta}(x_{t},t,s)$.
@@ 3.4 Unconditional Sampling Strategy
For unconditional sampling, we train an auxiliary latent diffusion model $p(s_{t-1}|s_{t},t)$ to sample $s$. During training, we first jointly optimize the prototypical encoder with the conditional DDPM; subsequently, the latent diffusion model is trained with the parameters of prototype encoder frozen.
@@ 3.5 Adding the Condition to the Objective
In Patronus, the prototype encoder is jointly optimized with the denoiser. To show that this simultaneous training does not degrade the generated distribution, we analyze how updating the condition $s$ affects the likelihood. For ease of derivation, we use the Evidence Lower Bound (ELBO) as an equivalent objective to generalize denoising losses (ho2020denoising).
@@
Any ELBO-improving update for the encoder always leads to progress. Such an update either increases the conditional likelihood of the data under the model or reduces the KL divergence between the generated distribution and the underlying data distribution.
@@
Let $z=x_{1:T}$ denote the forward noised latents generated from a fixed forward process $q(z\mid x)$, as is standard in diffusion models. For parameters $s$, define the per-sample evidence lower bound (ELBO)
$\mathrm{ELBO}(x;s)=\mathbb{E}_{q(z\mid x)}\!\left[\log p_{\theta}(x,z\mid s)-\log q(z\mid x)\right],$
and the conditional log-likelihood
$\log p_{\theta}(x\mid s)=\mathrm{ELBO}(x;s)+\mathrm{KL}\!\left(q(z\mid x)\,\big\|\,p_{\theta}(z\mid x,s)\right).$
Consider an update $s^{i}\mapsto s^{i+1}$ that satisfies $\mathrm{ELBO}(x;s^{i+1})\geq\mathrm{ELBO}(x;s^{i})$ for the data point $x$. Then the following identity holds:
$\log p_{\theta}(x\mid s^{i+1})-\log p_{\theta}(x\mid s^{i})=\big[\mathrm{ELBO}(x;s^{i+1})-\mathrm{ELBO}(x;s^{i})\big]+\Delta\mathrm{KL}(x),$ (7)
where $\Delta\mathrm{KL}(x)=\mathrm{KL}\!\left(q(z\mid x)\,\|\,p_{\theta}(z\mid x,s^{i+1})\right)-\mathrm{KL}\!\left(q(z\mid x)\,\|\,p_{\theta}(z\mid x,s^{i})\right).$
Given that the $\mathrm{ELBO}$ difference term is positive, Equation 7 implies that one of the following (improving) outcomes must occur:
Conditional likelihood improves. If $\Delta\mathrm{KL}(x)\leq 0$, then
$\mathrm{KL}\!\left(q(z\mid x)\,\|\,p_{\theta}(z\mid x,s^{i+1})\right)<=\mathrm{KL}\!\left(q(z\mid x)\,\|\,p_{\theta}(z\mid x,s^{i})\right).$
meaning that the model posterior $p_{\theta}(z\mid x,s^{i+1})$ more closely matches the forward distribution $q(z\mid x)$. This reduces the gap between the ELBO and the true conditional log-likelihood, yielding a tighter variational bound.
Generated samples become more probable. If $\Delta\mathrm{KL}(x)>0$ and considering that ELBO is also increased, then necessarily
$\log p_{\theta}(x\mid s^{i+1})\geq\log p_{\theta}(x\mid s^{i}),$
i.e. the model assigns higher probability to $x$.
Hence, any update of the encoder that increases the ELBO either increases the conditional likelihood of the data or reduces the KL divergence between the model posterior and the forward process, guaranteeing progress in approximating the data distribution. ∎
This analysis is inapplicable under alternative training objectives for latent condition (e.g., InfoDiff).